OpenStudy (anonymous):

3(y – 5) + 2 = 5 solve

OpenStudy (anonymous):

\[3(y-5)+2=5\]\[3y-15+2=5\]\[3y-13=5\]\[3y=18\]\[y=6\]

OpenStudy (anonymous):

Wait, where did you get 3+2=5 from?

OpenStudy (anonymous):

3(y-5)+2=5 first solve what's enclosed , thats 3(y-5) ... gives us 3y-15, now bring all the numbers on one side of the equation, so we have 3y-15+2=5 3y=5+15-2 3y=18 y=6